Solve the following differential equation by Laplace transforms. The function is subject to the given conditions. y +49y =
Answer #1

y +49y=0, yo) = 0, yło) = 1 using Laplace transform for egn- SYIS) – S YO) - YO) + 49 YCS) = 0 $ y(s) - 0-1 + 49 YS) = 0 {